OpenStudy (anonymous):

A Motorist travels xkm at 40km/h and y at 30km/h.The total time is 2 and a half hours.If the time taken to travel 6xkm at 30km/h and 4y at 50km/h 8s 14 hrs .Find x and y

ganeshie8 (ganeshie8):

time = distance / speed

ganeshie8 (ganeshie8):

A Motorist travels xkm at 40km/h and y at 30km/h.The total time is 2 and a half hours. \(\large \frac{x}{40} + \frac{y}{30} = 2.5\)

ganeshie8 (ganeshie8):

similarly setup another equation, and solve x and y ?

OpenStudy (anonymous):

\[t=\frac{d}{r} \]Solve the following for x and y:\[\left\{\frac{x}{40}+\frac{y}{30}=\frac{25}{10},\frac{6 x}{30}+\frac{4 y}{50}=14\right\} \]\[\left\{x=\frac{400 \text{ km}}{7},y=\frac{225 \text{ km}}{7}\right\} \]

OpenStudy (anonymous):

On an IMac, the solution should be viewed with the Firefox browser.